OpenStudy (anonymous):

solve following equation 9x-5-11x=5(3x-7)

OpenStudy (anonymous):

30/17

OpenStudy (anonymous):

9x-5-11x=5(3x-7)===> 9x-5-11x=15x-35 ===> -5+35=15x+11x-9x ===> 30=17x ===> x=30/17

OpenStudy (anonymous):

Simplifying 9x + -5 + -11x = 5(3x + -7) Reorder the terms: -5 + 9x + -11x = 5(3x + -7) Combine like terms: 9x + -11x = -2x -5 + -2x = 5(3x + -7) Reorder the terms: -5 + -2x = 5(-7 + 3x) -5 + -2x = (-7 * 5 + 3x * 5) -5 + -2x = (-35 + 15x) Solving -5 + -2x = -35 + 15x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-15x' to each side of the equation. -5 + -2x + -15x = -35 + 15x + -15x Combine like terms: -2x + -15x = -17x -5 + -17x = -35 + 15x + -15x Combine like terms: 15x + -15x = 0 -5 + -17x = -35 + 0 -5 + -17x = -35 Add '5' to each side of the equation. -5 + 5 + -17x = -35 + 5 Combine like terms: -5 + 5 = 0 0 + -17x = -35 + 5 -17x = -35 + 5 Combine like terms: -35 + 5 = -30 -17x = -30 x = -30/-17 = 30/17

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

i am frightened seeing the solution. is it such complicated?